OpenStudy (anonymous):

Question 17 The numbers in the circles go together in a certain way. Find the missing number marked by the question mark: 180, 14 120, 8 ?, 4 220,18

OpenStudy (anonymous):

there a pic with this problem?

ganeshie8 (ganeshie8):

80

OpenStudy (anonymous):

can you please explain how you got 80?

ganeshie8 (ganeshie8):

180 = (14+4) x 10 120 = (8 +4) x 10 220 = (18+4) x10 so... ? = (4+4) x 10 = 80